OpenStudy (anonymous):

Hexagon A is a regular hexagon. The total length of all the sides of the hexagon is 24 inches. Hexagon A is dilated about its center to create Hexagon B. The length of each side of Hexagon B is 4 and 4 over 5 inches. By what factor was Hexagon A dilated to create Hexagon B?

OpenStudy (anonymous):

5 over 6 1 and 1 over 5 2 4 over 5

OpenStudy (anonymous):

@experimentX

OpenStudy (experimentx):

first find the length of original hexagon, it is given by \( \frac{total\; length}{no \; of \;sides}\) then dilation factor = \( \frac{new\; length}{original \;length}\)

OpenStudy (anonymous):

i got @experimentX 4/5

OpenStudy (experimentx):

it shouldn't be ... what did you get for original length of hexagon??

OpenStudy (anonymous):

the original lenght is 4 isnt it

OpenStudy (experimentx):

oh ... (24/5)/4 = ??

OpenStudy (anonymous):

doesnt a hexagon have 6 sides ?

OpenStudy (anonymous):

@experimentX

OpenStudy (experimentx):

yes!!

OpenStudy (anonymous):

then wat did u get ? @experimentX

OpenStudy (experimentx):

what did you get first??

OpenStudy (anonymous):

4 over 5 @experimentX

OpenStudy (experimentx):

original length is 4 new length is 24/5 so you dilation factor is \( \frac{24/5}{4} = ??\) <-- this is what i got.

OpenStudy (anonymous):

oooooo

OpenStudy (experimentx):

so answer is ??

OpenStudy (anonymous):

1.2

OpenStudy (experimentx):

6/5 = 1 and 1 over 5 = 1.2

OpenStudy (anonymous):

thank you i get it